OpenStudy (anonymous):

Show all steps used to solve the problem. One point for steps and one point for solution. Be sure that you check your solution for correctness. If 3a - 9 = 6, what is the value of 4a - 1?

OpenStudy (anonymous):

given 3a -9=6 find 4a -1 3a= 6 +9 3a=15 3a/3=15/3 a=5 4a -1=? 4(5) -1 20 - 1 19